OpenStudy (anonymous):

Solve the following system of equations. -2x + 4y = 70 x + 5y = 28 @KamiBug

OpenStudy (kamibug):

Solve the second equation for the value of x. Then take that value and substitute it for x in the first equation so that you can solve it for y. Once you know what y is, substitute the value of y into the second equation to find x. x + 5y = 28 x = 28 - 5y -2x + 4y = 70 -2(28-5y) + 4y = 70 -56 + 10y + 4y = 70 14y = 126 y = 9 <--- x + 5(9) = 28 x + 45 = 28 x = ? Hope this helps! :)

OpenStudy (anonymous):

thank you

OpenStudy (kamibug):

Sure ^.^